If 4xy−kx2y−5y3 is the imaginary part of the analytic function f(z)=u+iv then, the value of k
1
Expert's answer
2020-04-30T20:23:56-0400

By the Cauchy–Riemann equations we have "\frac{\partial u}{\partial x}=\frac{\partial v}{\partial y}=4x-kx^2-15y^2" and "\frac{\partial u}{\partial y}=-\frac{\partial v}{\partial x}=-4y+2kxy"

Then "\frac{\partial^2 u}{\partial x\partial y}=\frac{\partial}{\partial x}\left(\frac{\partial u}{\partial y}\right)=2ky" and "\frac{\partial^2 u}{\partial x\partial y}=\frac{\partial}{\partial y}\left(\frac{\partial u}{\partial x}\right)=-30y". So "2k=-30" and "k=-15"

Answer: "k=-15"